# Stream: Scaling up Mechanistic Interpretability to Long Context in LLMs via Sparse Attention

### v1

@@ Abstract
As Large Language Models (LLMs) scale to million-token contexts, traditional Mechanistic Interpretability techniques for analyzing attention scale quadratically with context length, demanding terabytes of memory beyond 100,000 tokens. We introduce Sparse Tracing, a novel technique that leverages dynamic sparse attention to efficiently analyze long context attention patterns. We present Stream, a compilable hierarchical pruning algorithm that estimates per-head sparse attention masks in near-linear time $O(T\log T)$ and linear space $O(T)$, enabling one-pass interpretability at scale. Stream performs a binary-search-style refinement to retain only the top-$k$ key blocks per query while preserving the model’s next-token behavior. We apply Stream to long chain-of-thought reasoning traces and identify thought anchors while pruning 97-99% of token interactions. On the RULER benchmark, Stream preserves critical retrieval paths while discarding 90-96% of interactions and exposes layer-wise routes from the needle to output. Our method offers a practical drop-in tool for analyzing attention patterns and tracing information flow without terabytes of caches. By making long context interpretability feasible on consumer GPUs, Sparse Tracing helps democratize chain-of-thought monitoring. Code is available at https://anonymous.4open.science/r/stream-03B8/.
@@ 1 Introduction
Extending the context length of Large Language Models (LLMs) has become a major research focus with frontier models providing 1M token context length as standard [5, 38, 26, 10]. For complex tasks, longer contexts enable longer chain-of-thoughts in reasoning models which can improve performance [18]. In RAG applications, longer context lengths complement retrieval mechanisms, allowing models to consume larger numbers of documents at inference time [21].
Mechanistic interpretability seeks to reverse engineer neural networks such as LLMs, often by studying smaller toy models such as GPT-2, or decomposing larger models into simpler units that can be analyzed independently [27]. However, these common techniques such as those introduced by Lindsey et al. [22] and Nanda [27] face scaling challenges when applied to long context scenarios. First, meaningful signals become increasingly diffuse as they spread across extended contexts, making it difficult to isolate and detect interpretable patterns. Second, the attention mechanism’s quadratic scaling - $O(T^{2})$ in both computational time and memory usage - creates prohibitive resource demands, requiring terabytes of memory to cache all attention patterns for contexts with 100,000 tokens or more.
Scaling efforts in mechanistic interpretability have predominantly focused on accommodating larger model parameters, while the impact of extended context lengths remains significantly underexplored. Many recent studies explicitly defer interpretability analysis of contexts exceeding 100 tokens to future work [28, 22, 33], highlighting this critical gap in the field. This paper directly addresses these scalability challenges by introducing an interpretability technique that seamlessly scales to contexts of 100,000 tokens on consumer-grade GPUs, helping democratize long context mechanistic interpretability in realistic settings.
@@ 2 Background
The standard vanilla self-attention computation on current hardware accelerators is memory-bound, scaling with $O(T^{2})$ where $T$ is context length. For example, consider a standard LLM using bfloat16 quantization and context length of 100,000 tokens. A single attention pattern would require 20GB of VRAM to realize. Caching all attention patterns across the 34 layers and 8 heads in Gemma 3 4B [32] would require 3.84TB.
FlashAttention [11] is an exact attention algorithm that is computationally equivalent to vanilla attention but achieves $O(T)$ complexity by chunking the $T\times T$ attention pattern into blocks and computing them sequentially. Other approaches achieve sub-quadratic memory complexity by only computing a subset of each attention pattern, applying a sparse mask in three different ways: (1) the popular static sliding window patterns seen in LLMs designed for long context [5, 38, 32], (2) natively trainable sparse attention (NSA) [37] and (3) dynamically generated sparse attention at inference time [20, 39].
A lesser challenge LLMs face as context lengths increase is the quadratically scaled memory footprint of the Key-Value (KV) cache during inference. Grouped-Query Attention (GQA) [1] allows multiple query heads to share a single key and value head, significantly compressing the cache size with minimal impact on performance. As such, this is often not the bottleneck as context lengths increase. Additionally, long context LLMs are often pretrained on shorter contexts and Rotary Positional Embeddings (RoPE) [31] extensions such as YaRN [29], NTK-Aware interpolation [30] and position interpolation [9] are used to extend their context length with no impact on time and space complexity [41].
@@
Lee et al. [20] introduce Hierarchically Pruned (HiP) Attention to address the quadratic scaling challenges introduced by vanilla attention. The transformer’s self-attention layer computes a dense score matrix:
@@ 3 Method
In this paper we introduce Sparse Tracing, a technique comprising a set of mechanistic interpretability algorithms that leverage sparse attention to efficiently analyze attention patterns in long context scenarios. Sparse attention methods reduce computational complexity of the attention computation by computing only the most relevant portions of attention patterns. We hypothesize this may be a useful method of identifying parts of the model related to the model output. Additionally, we observe that the same computational bottlenecks that affect attention computation during inference also hinder interpretability analysis. Additionally, many mechanistic interpretability techniques scale linearly with the number of model components being explored, or require multiple forward and/or backwards passes [27, 19, 25]. Our technique explores all components at once, and requires only a single forward pass when the desired sparsity constant $k$ is known, and $\left\lfloor log_{2}k\right\rfloor$ forward passes otherwise, where the wall clock time of a single forward pass with our chosen Sparse Tracing algorithm is less than two forward passes without hooks.
@@
We introduce Stream, our specific Sparse Tracing algorithm based heavily on HiP attention [20]. Stream is visualized in Figure 1 and the algorithm is provided in full in Appendix C. HiP attention is a compilable, dynamic sparse attention algorithm that achieves $O(T\log T)$ time complexity and $O(T)$ space complexity and allows the granularity of the attention computation to be tuned. Stream is heavily based on HiP attention and operates via the same hierarchical search process that efficiently identifies the most relevant blocks of keys for each block of queries.
@@
Our criteria of $n_{\text{match}}=2$ matching tokens generated in a row is a proxy for perplexity and model performance. We ablate this choice in Appendix A.2. Lee et al. [20] advise against replacing the first few decoder layers $l_{d}$ of the transformer with their sparse HiP attention as they found that the first few layers have substantially denser attention patterns and pruning these can degrade performance severely. Through ablation studies they recommend $l_{d}=3$ which we ablate in Appendix A.1.
@@ 4.1 Case Study 1: Thought Anchors
Bogdan et al. [6] introduce black-box and white-box approaches to interpreting long context chain-of-thought reasoning traces by aggregating attention patterns between pairs of sentences rather than between pairs of tokens. They coin the term “thought anchors”, which are steps in the reasoning trace that have a disproportionately large influence on the remainder of the trace. In this case study, we validate Sparse Tracing as a highly scalable method - in both time and space complexity - of identifying thought anchors.
@@
We follow the experimental setup of Bogdan et al. [6]. Stream can be applied to any LLM that implements attention, and for this case study we report results on DeepSeek R1-Distill Qwen-1.5B [14, 36] a small reasoning in model in the same family as Bogdan et al. [6]’s prior work. Matching Bogdan et al. [6], we use a temperature of 0.6 and a top-p value of 0.95 to generate responses to problems in the MATH-500 dataset [15]. In a similar manner to Bogdan et al. [6], we use the results of running the model on 500 problems, 10 times each to identify 10 questions that the model correctly solves 25-75% of the time. Venhoff et al. [34] identifies a selection of reasoning behaviors exhibited by DeepSeek-R1-Distill models that can be controlled using steering vectors. Based on this framework, Bogdan et al. [6] prompt an LLM to label sentences as one of 8 distinct categories. These categories are (1) problem setup, (2) plan generation, (3) fact retrieval, (4) active computation, (5) uncertainty management, (6) result consolidation, (7) self-checking and (8) final answer emission. In our experiments, instead of splitting the reasoning trace into sentences, we split the traces into 32 token chunks to match the query block size ($b_{q}$) in Stream. We prompt OpenAI GPT-4o (April-May, 2025) [17] to label each sentence using an almost identical prompt given by Bogdan et al. [6] which we provide in Appendix D.1.
@@ 4.1.1 Receiver Heads
Bogdan et al. [6] suggest that “thought anchors” attract higher attention from later parts of the text. In Sparse Tracing, we adopt a similar idea: salient query blocks should focus on a small, highly relevant set of key blocks. Formally, let
@@
That is, more salient queries appear more often among top-$k$ pairs. To detect such anchors, we look for attention heads with high kurtosis, meaning a few parts of the context receive much stronger attention than the rest. Following Bogdan et al. [6], we call these receiver heads. They concentrate attention on specific blocks instead of spreading it broadly. Instead of aggregating attention to the sentence level, we compute block-level averages over $b_{q}\times b_{k}$ spans, yielding the “block mean” pattern (Figure 2). We select potential receiver heads by measuring the maximum vertical kurtosis of these block means at multiple context lengths. For each candidate head, we generate sparse masks with Stream, both binary and score-weighted, and visualize them in Figure 2.
Figure 2: This figure displays attention patterns for three different context lengths (3,000, 6000, and 10000 tokens) using three different approaches: Full Attention Pattern (first column), Mean Averaged [6] with $b=32$ (second column), and Stream with $b_{y}=b_{k}=32$ (third and fourth column).
@@ 4.1.2 Sentence Category Importance
In Figure 3, we compare the sentence category importance based on vertical attention scores, comparing Bogdan et al. [6]’s approach and Stream.
@@ 4.2 Case Study 2: Needle in a Haystack
In this case study, we explore the scalability of Stream by applying it to the RULER needle in a haystack benchmark [16]. Stream can be applied to any LLM that implements attention, and we validate the transferability of our method by applying it to a different LLM, namely Gemma 3 1B [32].
@@
The experimental data was prepared using the RULER benchmark’s [16] synthetic data generation pipeline, which creates structured evaluation samples with controllable context lengths and needle placement positions. For each evaluation instance, we employed a binary search algorithm to determine the optimal sparsity parameter $k$ - the minimum number of key-value blocks required for the sparse attention mechanism to successfully retrieve two consecutive tokens from the needle (see Appendix A.2 for ablations).
@@ 4.2.1 Sparse Tracing Needles
Figure 4: Visualization of sparse attention masks generated by Stream on the RULER needle-in-a-haystack benchmark [16] at three context lengths (1k, 10k, and 20k tokens). The needle is marked using a red vertical line in the Attention Pattern plots.
@@ 4.2.2 Exploring the effect of Needle Depth
In needle-in-a-haystack tasks, models often show a U-shaped retrieval curve [23]: they are relatively good at retrieving needles near the start of the context (due to over-squashing) and near the end (as those positions are directly learned), but struggle in the middle. By introducing the effective sparsity $s$ such that:
@@ 4.2.3 Exploring Information Flow
Existing methods for tracing information flow in transformers [13, 27, 19, 25] scale quadratically with context length. To our knowledge, Stream is the first such technique with linear scaling, making it feasible for tasks like indirect object identification and needle-in-a-haystack. For the latter, we report sparse patterns at $T=1000$, needle depth $n_{d}\approx 0.5$, $b_{q}=b_{k}=32$, with $k_{\text{success}}=6$ (successful retrieval) and $k_{\text{fail}}=3$ (unsuccessful retrieval). Subtracting the $k_{\text{fail}}$ masks from $k_{\text{success}}$ across layers and heads isolates the connections that enable retrieval. Figure 6 visualizes these paths, weighted by block scores: red edges trace needle-to-output flows, while blue edges denote other attention routes. We restrict analysis to attention paths, leaving residual and MLP components for future Sparse Tracing extensions. Full patterns for $k_{\text{success}}$ and $k_{\text{fail}}$ are shown in Appendix E.1.
@@ Discussion.
We see that the needle information appears to travel along two paths to the output, through hidden state blocks 18 and 19, and through the final few blocks. We also see attention sink behavior - high attention to earlier tokens in the context - when needle information passes to the output block, which may be preventing over-mixing, a form of rank collapse [4]. This would aid in preserving the information from the needle through the layers of the network. This experiment consolidates the use of Stream as an empirical tool to explore theoretical claims.
@@ 5 Related Work
Reasoning models produce a long chain of thought before responding to the user, with the aim of increasing their performance on complex tasks [18]. A growing branch of Mechanistic Interpretability research seeks to understand model internals during chain-of-thought processes [6, 8, 7, 12, 34]. Brinkmann et al. [7] demonstrate that transformers can implement interpretable backward chaining algorithms for pathfinding, while Dutta et al. [12] extend this understanding to production-scale models, revealing multiple parallel pathways for answer generation. Venhoff et al. [34] show these reasoning behaviors can be controlled through steering vectors. Cabannes et al. [8] find identifiable computational structures that enable iterative algorithms, and Bogdan et al. [6] take black-box and white-box approaches to identify "thought anchors" - the most critical parts of the reasoning trace. In our work, we use our Sparse Tracing methodology to scalably identify "thought anchors" in longer contexts.
@@
Other work examines how LLMs process extended contexts and the mechanisms underlying their successes and failures [35, 23, 3, 2, 24]. Wu et al. [35] identify "retrieval heads" - a sparse set of attention heads (3-6%) specifically responsible for copying information from arbitrary locations in long contexts, providing a mechanistic explanation for factual retrieval capabilities. This work builds on Liu et al. [23]’s empirical finding that models perform poorly when relevant information appears in the middle of long contexts. Barbero et al. [2] provide theoretical grounding for these limitations, proving that decoder-only transformers suffer from "information over-squashing" where distinct input contexts become indistinguishable in final token representations. Their companion work [3] reveals how rotary positional embeddings can mitigate some issues by using different frequency components for positional versus semantic attention. Our work introduces Stream, which builds highly scalable computation graphs for information propagation research into Geometric Deep Learning metrics such as over-smoothing, over-squashing and under-reaching.
@@
Tooling has been built for conducting Mechanistic Interpretability research as model size, complexity and context lengths scale [27, 19, 13, 40]. Nanda [27] introduces attribution patching, enabling activation patching experiments on large models through gradient-based approximations. Kramár et al. [19] address failure modes in this approach, proposing AtP*. Ferrando and Voita [13] develop automated methods for discovering information flow routes without requiring human-designed counterfactuals, while Zheng et al. [40] provide systematic frameworks for understanding attention head functions. Paulo et al. [28] address the scalability challenge of interpreting sparse autoencoder (SAE) features by developing an automated pipeline that can generate and evaluate natural language explanations for millions of features. Templeton et al. [33] successfully scale SAEs to Claude 3 Sonnet extracting millions of interpretable features. We introduce the first Mechanistic Interpretability tool designed to run on consumer hardware that scales to long context.

### v2

@@ Abstract
As Large Language Models (LLMs) scale to million-token contexts, traditional Mechanistic Interpretability techniques for analyzing attention scale quadratically with context length, demanding terabytes of memory beyond 100,000 tokens. We introduce Sparse Tracing, a novel technique that leverages dynamic sparse attention to efficiently analyze long context attention patterns. We present Stream, a compilable hierarchical pruning algorithm that estimates per-head sparse attention masks in near-linear time $O(T\log T)$ and linear space $O(T)$, enabling one-pass interpretability at scale. Stream performs a binary-search-style refinement to retain only the top-$k$ key blocks per query while preserving the model’s next-token behavior. We apply Stream to long chain-of-thought reasoning traces and identify thought anchors while pruning 97-99% of token interactions. On the RULER benchmark, Stream preserves critical retrieval paths while discarding 90-96% of interactions and exposes layer-wise routes from the needle to output. Our method offers a practical drop-in tool for analyzing attention patterns and tracing information flow without terabytes of caches. By making long context interpretability feasible on consumer GPUs, Sparse Tracing helps democratize chain-of-thought monitoring. Code is available at https://github.com/spotify-research/stream-mechinterp/.
@@ 1 Introduction
Extending the context length of Large Language Models (LLMs) has become a major research focus with frontier models providing 1M token context length as standard [beltagy2020longformer, zaheer2020big, meta2025llama, comanici2025gemini]. For complex tasks, longer contexts enable longer chain-of-thoughts in reasoning models which can improve performance [jaech2024openai]. In RAG applications, longer context lengths complement retrieval mechanisms, allowing models to consume larger numbers of documents at inference time [li2024retrieval].
Mechanistic interpretability seeks to reverse engineer neural networks such as LLMs, often by studying smaller toy models such as GPT-2, or decomposing larger models into simpler units that can be analyzed independently [nanda2023attribution]. However, these common techniques such as those introduced by lindsey2025biology and nanda2023attribution face scaling challenges when applied to long context scenarios. First, meaningful signals become increasingly diffuse as they spread across extended contexts, making it difficult to isolate and detect interpretable patterns.
Second, the attention mechanism’s quadratic scaling - $O(T^{2})$ in both computational time and memory usage - creates prohibitive resource demands, requiring terabytes of memory to cache all attention patterns for contexts with 100,000 tokens or more.
Scaling efforts in mechanistic interpretability have predominantly focused on accommodating larger model parameters, while the impact of extended context lengths remains significantly underexplored. Many recent studies explicitly defer interpretability analysis of contexts exceeding 100 tokens to future work [paulo2024automatically, lindsey2025biology, templeton2024scaling], highlighting this critical gap in the field. This paper directly addresses these scalability challenges by introducing an interpretability technique that seamlessly scales to contexts of 100,000 tokens on consumer-grade GPUs, helping democratize long context mechanistic interpretability in realistic settings.
@@ 2 Background
The standard vanilla self-attention computation on current hardware accelerators is memory-bound, scaling with $O(T^{2})$ where $T$ is context length. For example, consider a standard LLM using bfloat16 quantization and context length of 100,000 tokens. A single attention pattern would require 20GB of VRAM to realize. Caching all attention patterns across the 34 layers and 8 heads in Gemma 3 4B [team2025gemma] would require 3.84TB.
FlashAttention [dao2022flashattention] is an exact attention algorithm that is computationally equivalent to vanilla attention but achieves $O(T)$ complexity by chunking the $T\times T$ attention pattern into blocks and computing them sequentially. Other approaches achieve sub-quadratic memory complexity by only computing a subset of each attention pattern, applying a sparse mask in three different ways: (1) the popular static sliding window patterns seen in LLMs designed for long context [beltagy2020longformer, zaheer2020big, team2025gemma], (2) natively trainable sparse attention (NSA) [yuan2025native] and (3) dynamically generated sparse attention at inference time [lee2024hip, zhang2025spargeattn].
A lesser challenge LLMs face as context lengths increase is the quadratically scaled memory footprint of the Key-Value (KV) cache during inference. Grouped-Query Attention (GQA) [ainslie2023gqa] allows multiple query heads to share a single key and value head, significantly compressing the cache size with minimal impact on performance. As such, this is often not the bottleneck as context lengths increase. Additionally, long context LLMs are often pretrained on shorter contexts and Rotary Positional Embeddings (RoPE) [su2024roformer] extensions such as YaRN [peng2023yarn], NTK-Aware interpolation [roziere2023code] and position interpolation [chen2023extending] are used to extend their context length with no impact on time and space complexity [zhong2024understanding].
@@
lee2024hip introduce Hierarchically Pruned (HiP) Attention to address the quadratic scaling challenges introduced by vanilla attention. The transformer’s self-attention layer computes a dense score matrix:
@@ 3 Method
In this paper we introduce Sparse Tracing, a technique comprising a set of mechanistic interpretability algorithms that leverage sparse attention to efficiently analyze attention patterns in long context scenarios. Sparse attention methods reduce computational complexity of the attention computation by computing only the most relevant portions of attention patterns. We hypothesize this may be a useful method of identifying parts of the model related to the model output. Additionally, we observe that the same computational bottlenecks that affect attention computation during inference also hinder interpretability analysis. Additionally, many mechanistic interpretability techniques scale linearly with the number of model components being explored, or require multiple forward and/or backwards passes [nanda2023attribution, kramar2024atp, meng2022locating]. Our technique explores all components at once, and requires only a single forward pass when the desired sparsity constant $k$ is known, and $\left\lfloor log_{2}k\right\rfloor$ forward passes otherwise, where the wall clock time of a single forward pass with our chosen Sparse Tracing algorithm is less than two forward passes without hooks.
@@
We introduce Stream, our specific Sparse Tracing algorithm based heavily on HiP attention [lee2024hip]. Stream is visualized in Figure 1 and the algorithm is provided in full in Appendix C. HiP attention is a compilable, dynamic sparse attention algorithm that achieves $O(T\log T)$ time complexity and $O(T)$ space complexity and allows the granularity of the attention computation to be tuned. Stream is heavily based on HiP attention and operates via the same hierarchical search process that efficiently identifies the most relevant blocks of keys for each block of queries.
@@
Our criteria of $n_{\text{match}}=2$ matching tokens generated in a row is a proxy for perplexity and model performance. We ablate this choice in Appendix A.2. lee2024hip advise against replacing the first few decoder layers $l_{d}$ of the transformer with their sparse HiP attention as they found that the first few layers have substantially denser attention patterns and pruning these can degrade performance severely. Through ablation studies they recommend $l_{d}=3$ which we ablate in Appendix A.1.
@@ 4.1 Case Study 1: Thought Anchors
bogdan2025thought introduce black-box and white-box approaches to interpreting long context chain-of-thought reasoning traces by aggregating attention patterns between pairs of sentences rather than between pairs of tokens. They coin the term “thought anchors”, which are steps in the reasoning trace that have a disproportionately large influence on the remainder of the trace. In this case study, we validate Sparse Tracing as a highly scalable method - in both time and space complexity - of identifying thought anchors.
@@
We follow the experimental setup of bogdan2025thought. Stream can be applied to any LLM that implements attention, and for this case study we report results on DeepSeek R1-Distill Qwen-1.5B [guo2025deepseek, yang2024qwen2] a small reasoning in model in the same family as bogdan2025thought’s prior work. Matching bogdan2025thought, we use a temperature of 0.6 and a top-p value of 0.95 to generate responses to problems in the MATH-500 dataset [hendrycks2021measuring]. In a similar manner to bogdan2025thought, we use the results of running the model on 500 problems, 10 times each to identify 10 questions that the model correctly solves 25-75% of the time. venhoff2025understanding identifies a selection of reasoning behaviors exhibited by DeepSeek-R1-Distill models that can be controlled using steering vectors. Based on this framework, bogdan2025thought prompt an LLM to label sentences as one of 8 distinct categories. These categories are (1) problem setup, (2) plan generation, (3) fact retrieval, (4) active computation, (5) uncertainty management, (6) result consolidation, (7) self-checking and (8) final answer emission. In our experiments, instead of splitting the reasoning trace into sentences, we split the traces into 32 token chunks to match the query block size ($b_{q}$) in Stream. We prompt OpenAI GPT-4o (April-May, 2025) [hurst2024gpt] to label each sentence using an almost identical prompt given by bogdan2025thought which we provide in Appendix D.1.
@@ 4.1.1 Receiver Heads
bogdan2025thought suggest that “thought anchors” attract higher attention from later parts of the text. In Sparse Tracing, we adopt a similar idea: salient query blocks should focus on a small, highly relevant set of key blocks. Formally, let
@@
That is, more salient queries appear more often among top-$k$ pairs. To detect such anchors, we look for attention heads with high kurtosis, meaning a few parts of the context receive much stronger attention than the rest. Following bogdan2025thought, we call these receiver heads. They concentrate attention on specific blocks instead of spreading it broadly. Instead of aggregating attention to the sentence level, we compute block-level averages over $b_{q}\times b_{k}$ spans, yielding the “block mean” pattern (Figure 2). We select potential receiver heads by measuring the maximum vertical kurtosis of these block means at multiple context lengths. For each candidate head, we generate sparse masks with Stream, both binary and score-weighted, and visualize them in Figure 2.
Figure 2: This figure displays attention patterns for three different context lengths (3,000, 6000, and 10000 tokens) using three different approaches: Full Attention Pattern (first column), Mean Averaged [bogdan2025thought] with $b=32$ (second column), and Stream with $b_{y}=b_{k}=32$ (third and fourth column).
@@ 4.1.2 Sentence Category Importance
In Figure 3, we compare the sentence category importance based on vertical attention scores, comparing bogdan2025thought’s approach and Stream.
@@ 4.2 Case Study 2: Needle in a Haystack
In this case study, we explore the scalability of Stream by applying it to the RULER needle in a haystack benchmark [hsieh2024ruler]. Stream can be applied to any LLM that implements attention, and we validate the transferability of our method by applying it to a different LLM, namely Gemma 3 1B [team2025gemma].
@@
The experimental data was prepared using the RULER benchmark’s [hsieh2024ruler] synthetic data generation pipeline, which creates structured evaluation samples with controllable context lengths and needle placement positions. For each evaluation instance, we employed a binary search algorithm to determine the optimal sparsity parameter $k$ - the minimum number of key-value blocks required for the sparse attention mechanism to successfully retrieve two consecutive tokens from the needle (see Appendix A.2 for ablations).
@@ 4.2.1 Sparse Tracing Needles
Figure 4: Visualization of sparse attention masks generated by Stream on the RULER needle-in-a-haystack benchmark [hsieh2024ruler] at three context lengths (1k, 10k, and 20k tokens). The needle is marked using a red vertical line in the Attention Pattern plots.
@@ 4.2.2 Exploring the effect of Needle Depth
In needle-in-a-haystack tasks, models often show a U-shaped retrieval curve [liu2023lost]: they are relatively good at retrieving needles near the start of the context (due to over-squashing) and near the end (as those positions are directly learned), but struggle in the middle. By introducing the effective sparsity $s$ such that:
@@ 4.2.3 Exploring Information Flow
Existing methods for tracing information flow in transformers [ferrando2024information, nanda2023attribution, kramar2024atp, meng2022locating] scale quadratically with context length. To our knowledge, Stream is the first such technique with linear scaling, making it feasible for tasks like indirect object identification and needle-in-a-haystack. For the latter, we report sparse patterns at $T=1000$, needle depth $n_{d}\approx 0.5$, $b_{q}=b_{k}=32$, with $k_{\text{success}}=6$ (successful retrieval) and $k_{\text{fail}}=3$ (unsuccessful retrieval). Subtracting the $k_{\text{fail}}$ masks from $k_{\text{success}}$ across layers and heads isolates the connections that enable retrieval. Figure 6 visualizes these paths, weighted by block scores: red edges trace needle-to-output flows, while blue edges denote other attention routes. We restrict analysis to attention paths, leaving residual and MLP components for future Sparse Tracing extensions. Full patterns for $k_{\text{success}}$ and $k_{\text{fail}}$ are shown in Appendix E.1.
@@ Discussion.
We see that the needle information appears to travel along two paths to the output, through hidden state blocks 18 and 19, and through the final few blocks. We also see attention sink behavior - high attention to earlier tokens in the context - when needle information passes to the output block, which may be preventing over-mixing, a form of rank collapse [barbero2025llms]. This would aid in preserving the information from the needle through the layers of the network. This experiment consolidates the use of Stream as an empirical tool to explore theoretical claims.
@@ 5 Related Work
Reasoning models produce a long chain of thought before responding to the user, with the aim of increasing their performance on complex tasks [jaech2024openai]. A growing branch of Mechanistic Interpretability research seeks to understand model internals during chain-of-thought processes [bogdan2025thought, cabannes2024iteration, brinkmann2024mechanistic, dutta2024think, venhoff2025understanding]. brinkmann2024mechanistic demonstrate that transformers can implement interpretable backward chaining algorithms for pathfinding, while dutta2024think extend this understanding to production-scale models, revealing multiple parallel pathways for answer generation. venhoff2025understanding show these reasoning behaviors can be controlled through steering vectors. cabannes2024iteration find identifiable computational structures that enable iterative algorithms, and bogdan2025thought take black-box and white-box approaches to identify "thought anchors" - the most critical parts of the reasoning trace. In our work, we use our Sparse Tracing methodology to scalably identify "thought anchors" in longer contexts.
@@
Other work examines how LLMs process extended contexts and the mechanisms underlying their successes and failures [wu2024retrieval, liu2023lost, barbero2024round, barbero2024transformers, liu2025survey]. wu2024retrieval identify "retrieval heads" - a sparse set of attention heads (3-6%) specifically responsible for copying information from arbitrary locations in long contexts, providing a mechanistic explanation for factual retrieval capabilities. This work builds on liu2023lost’s empirical finding that models perform poorly when relevant information appears in the middle of long contexts. barbero2024transformers provide theoretical grounding for these limitations, proving that decoder-only transformers suffer from "information over-squashing" where distinct input contexts become indistinguishable in final token representations. Their companion work [barbero2024round] reveals how rotary positional embeddings can mitigate some issues by using different frequency components for positional versus semantic attention. Our work introduces Stream, which builds highly scalable computation graphs for information propagation research into Geometric Deep Learning metrics such as over-smoothing, over-squashing and under-reaching.
@@
Tooling has been built for conducting Mechanistic Interpretability research as model size, complexity and context lengths scale [nanda2023attribution, kramar2024atp, ferrando2024information, zheng2024attention]. nanda2023attribution introduces attribution patching, enabling activation patching experiments on large models through gradient-based approximations. kramar2024atp address failure modes in this approach, proposing AtP*. ferrando2024information develop automated methods for discovering information flow routes without requiring human-designed counterfactuals, while zheng2024attention provide systematic frameworks for understanding attention head functions. paulo2024automatically address the scalability challenge of interpreting sparse autoencoder (SAE) features by developing an automated pipeline that can generate and evaluate natural language explanations for millions of features. templeton2024scaling successfully scale SAEs to Claude 3 Sonnet extracting millions of interpretable features. We introduce the first Mechanistic Interpretability tool designed to run on consumer hardware that scales to long context.